OpenStudy (beccaboo333):

What would the coordinates be? I'm not good with algebra. Explaining or helping me is greatly appreciated. :)

OpenStudy (beccaboo333):

x - 3y = -12 2x - y = 1

OpenStudy (johnweldon1993):

So I would use the elimination method here....have you learned this?

OpenStudy (beccaboo333):

... i don't understand algebra.. one bit.. only simple things. i honestly do not know how to do most of these things.

OpenStudy (texaschic101):

x - 3y = -12 ---> x = 3y - 12 now sub 3y - 12 in for x in the other equation 2x - y = 1 2(3y - 12) - y = 1 -- distribute through the parenthesis 6y - 24 - y = 1 -- combine like terms 5y - 24 = 1 -- add 24 to both sides 5y = 1 + 24 5y = 25 -- divide both sides by 5 y = 5 now sub 5 in for y in either of the original equations x - 3y = -12 x - 3(5) = -12 x - 15 = -12 x = -12 + 15 x = 3 check... 2x - y = 1 2(3) - 5 = 1 6 - 5 = 1 1 = 1 (correct) x = 3 and y = 5 any questions ?

OpenStudy (beccaboo333):

whaaaaaaaaaaaa?

OpenStudy (texaschic101):

I used the substitution method instead of the elimination method. I find it much easier. I just solved for a variable at the beginning, which was x = 3y - 12. I then subbed 3y - 12 in for x in the other equation to find y. Once I found y, I subbed that back into an equation and found x. I then check my answers by subbing them into an original equation, and they checked out

OpenStudy (beccaboo333):

Thank you. I sorta understand now.

OpenStudy (texaschic101):

so basically your coordinates are (3,5) Glad I could help :)